QUESTION 4

The mean number of cars confiscated by a wrecker is five per day. Using the

Poisson distribution,

a) find the probability of the number of cars confiscated on any given day is

exactly 5

b) At most 3

c) More than 3
Answer #1

X~ pay is pere day ) - X= 1) = e- کم کرد) کی Tecx=x+ 10.1755) b) P X < 3) = 3 E xe O. & 6so - PLX > 3) = 1 L- PCX <3 ) 3 I- z